Using the chemical reaction, Xe(g)+F 2 (g) XeF 4 (g) , how many ml of xexor tetrafluoride will be produced from 2.65 mL of xenon and 1.74 mL of fluorine?


1
Expert's answer
2022-04-14T11:59:04-0400

Xe(g) + 2F2(g) = XeF4(g)

n = V / 22.4

n (Xe) = 0.00265 / 22.4 = 0.0001 mol

n (F2) = 0.00174 / 22.4 = 0.00008 mol

According to the equation, n (XeF4) = 1/2 x n (F2) = n (Xe)

F2 is the limiting reagent.

n (XeF4) = 1/2 x n (F2) = 1/2 x 0.00008 = 0.00004 mol

V (XeF4) = 0.00004 x 22.4 = 0.00087 L = 0.87 ml